# Approximation in Hölder spaces

### v3

@@ Abstract.
We introduce new vanishing subspaces of the homogeneous Hölder space $\dot{C}^{0,\omega}(X,Y),$ in the generality of a doubling modulus $\omega$ and normed spaces $X$ and $Y.$ For many couples $X,Y,$ we show these vanishing subspaces to completely characterize those Hölder functions that admit approximations, in the Hölder seminorm, by smooth, Lipschitz and boundedly supported functions. Beyond the intrinsic interest of these results, we also present connections to bi-parameter harmonic analysis on the Euclidean space and in particular to compactness of the bi-commutator.
@@ 1.1. Introduction
$|||f|||_{C^{0,\omega}(X,Y)}=\|f\|_{\dot{C}^{0,\omega}(X,Y)}+\|f(0)\|_{Y}.$ (1)
Dropping the term $\|f(0)\|_{Y},$ the seminorm $\|\cdot\|_{\dot{C}^{0,\omega}(X,Y)}$ becomes a norm when the functions are identified with equivalence classes modulo additive constants. In the particular case $\omega(t)=t^{\alpha},$ $\alpha\in(0,1),$ we recover the Hölder spaces $\dot{C}^{0,\alpha}(X,Y).$ Moreover, provided $\omega$ is any reasonable continuity moduli – at least $\omega(t)\to 0,$ as $t\to 0$ – the spaces $\dot{C}^{0,\omega}(X,Y)$ become subspaces of uniformly continuous functions.
@@
In the present article, we prove full characterizations of approximation with Hölder type rate of convergence taking the place of uniform convergence. We note that these two types of convergence do not follow from each other; see Appendix A for a detailed comparison with simple observations and examples. Our main result is best described as the recognition of the vanishing scales small, large and far (below) as precisely the concept that allows a complete description of the closure of smooth, Lipschitz and boundedly supported functions with respect to the $\dot{C}^{0,\omega}$ seminorm. We notate
@@
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{small}}(X,Y)$ =$=\big{\{}f:X\to Y\,:\,\lim_{\delta\to 0}\operatorname{osc}^{\omega}_{\delta}(f)=0\big{\}},\,$
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{large}}(X,Y)$ =$=\big{\{}f:X\to Y\,:\,\lim_{\delta\to\infty}\operatorname{osc}^{\omega}_{\delta}(f)=0\big{\}},$
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{far}}(X,Y)$ =$=\big{\{}f:X\to Y\,:\,\lim_{\delta\to\infty}\sup_{\min(\|x\|,\|y\|)>\delta}\operatorname{osc}^{\omega}_{(x,y)}(f)=0\big{\}}.$
@@
For each scale $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ set
@@
If $Y$ is a Banach space, then so are all of $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ as spaces of modulo additive constant equivalence classes; for full details, see Remark 2.1 below. From now on, whenever no confusion is possible, we start dropping the source and target spaces from the norms and simplify notation with $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|;$ in the same way, we write $\dot{C}^{0,\omega}(X,Y)=\dot{C}^{0,\omega}(X)=\dot{C}^{0,\omega},$ and so on.
@@
where the doubling constant $C_{\operatorname{db}}>1.$ This is a big class of moduli that in particular covers the Hölder modulus $\alpha(t):=t^{\alpha},$ when $\alpha\in(0,1),$ but of course other more exotic moduli are also covered. The classes $\dot{\operatorname{VC}}^{0,\alpha}_{\operatorname{small}}$ are also called the little Hölder spaces. These are important in the study of Lipschitz algebras in metric spaces, and we refer to the monograph by Weaver [Weaverbook, Chapters 4 and 8] for a background. See also the recent monograph by D. Mitrea, I. Mitrea, and M. Mitrea [Mitreabook, Chapter 3] for a detailed exposition of the Hölder spaces in the setting of Ahlfors regular sets and their connection with functions of vanishing mean oscillation.
@@ 1.3. Applications on the Euclidean space
In the previous section we presented results that guarantee nice approximations, provided we know pointwise vanishing type conditions. When $X=\mathbb{R}^{n},$ we present formulations of these results in terms of mean oscillations type conditions, that are very convenient when studying the boundedness and compactness of commutators of Calderón-Zygmund operators (CZOs). In the one parameter setting our results recover important components to the theory, but from a more general approach. In the bi-parameter setting our results have very recently found applications by Martikainen and the second named author [MarOik24] in the study of the compactness of the bi-commutator. We describe all of this in full detail below, but to not get too far ahead of ourselves, let us begin by recapping boundedness and compactness characterizations of commutators in the one-parameter setting.
For a non-degenerate CZO $T$ on $\mathbb{R}^{n},$ a function $b\in L^{1}_{\operatorname*{loc}}(\mathbb{R}^{n},\mathbb{C})$ (symbol of the commutator) and exponents $p,q\in(1,\infty),$ the commutator has the following mapping properties
@@
where the space $X^{p,q}(\mathbb{R}^{n})$ is determined as follows. Given $p,q\in(1,\infty),$ denote by $\alpha(p,q)$ and $r(p,q)$ the exponents uniquely determined through the relations $\alpha(p,q)/n=1/p-1/q$ and $1/q=1/r(p,q)+1/p.$ Then, $X^{p,q}(\mathbb{R}^{n})$ is the space of bounded mean oscillations $\mathrm{BMO}(\mathbb{R}^{n}),$ when $p=q;$ the fractional BMO space $\mathrm{BMO}^{\alpha(p,q)}(\mathbb{R}^{n}),$ when $p<q;$ and the Lebesgue space $L^{r(p,q)}(\mathbb{R}^{n})$ modulo additive constants $\dot{L}^{r(p,q)}(\mathbb{R}^{n}),$ when $p>q.$ (The reader can recover the definitions of the spaces $\mathrm{BMO},\mathrm{BMO}^{\alpha}$ from Definition 1.7 below with the ``$\mbox{moduli}"$ $t\mapsto t^{0}$ and $t\mapsto t^{\alpha},$ respectively.) Emphasizing the recognition of the correct function spaces $X^{p,q}(\mathbb{R}^{n})$ to which the symbol belongs, the result (5) is due to Nehari [Nehari1957] ($n=1$) and Coifman, Rochberg, Weiss [CRW] ($n\geq 2$), when $q=p;$ due to Janson [Jan1978], when $q>p;$ and due to Hytönen [HyLpLq], when $q<p.$ We also refer the reader to the introductory section of [HyLpLq] for precise definitions and a fuller history of what was outlined above. Next, we describe the corresponding off-diagonal compactness characterization.
When $\alpha(p,q)>1,$ $\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n})$ consists of the constant functions and the commutator is bounded iff it is the zero operator (hence compact, in particular). Thus: for compactness only those exponents $p,q$ that result in $\alpha(p,q)\leq 1$ are interesting to study. When $\alpha(p,q)=1$ it is the case [GHWY21, Theorem 1.7. & 1.8.] that the commutator is compact iff it is the zero operator, again. Now for all $p,q\in(1,\infty)$ such that $\alpha(p,q)<1,$ there holds that
@@
The characterization (6) is due to Uchiyama [Uch1978], when $p=q$; due to Guo, He, Wu and Yang [GHWY21], when $q>p;$ and due to Hytönen, Li, Tao and Yang [HLTY2023], when $q<p.$ When $q>p$ and $\alpha(p,q)\in(0,1),$ a result of Meyer’s [Mey1964] states that $\mathrm{BMO}^{\alpha(p,q)}(\mathbb{R}^{n})=\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n}).$ Thus, the following Theorem 1.6 provides, in particular, a new characterization of the space $Y^{p,q}(\mathbb{R}^{n})$ in (6).
@@
To obtain the implications ``"``"`$\Rightarrow$"` ` ⇒ " in (6) it is very useful to have a description of $Y^{p,q}(\mathbb{R}^{n}),$ in the cases $q\geq p,$ in terms of some vanishing mean oscillation (VMO) criteria. Such VMO criteria (Definition 1.8, below) usually follow immediately from the existence of approximations, but to establish approximations from VMO criteria is delicate. Recall that in the previous section we established existence of approximations beginning from some pointwise vanishing criteria $\dot{\operatorname{VC}}.$ Below, as Theorem 1.9, we provide the correspondence between these $\mathrm{VMO}$ and the $\dot{\operatorname{VC}}$ type criterions, which allows us, when $X=\mathbb{R}^{n},$ to reformulate all of our results, among them Theorem 1.6, in terms of mean oscillations.
@@ 1.4. Formulation as mean oscillation conditions
$\mathrm{VMO}^{\omega}_{\operatorname{small}}(\mathbb{R}^{n},Y)$ =$=\Big{\{}f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to 0}\sup_{\begin{subarray}{c}\ell(Q)=\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big{\}},$
$\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)$ =$=\Big{\{}f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}\ell(Q)=\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big{\}},$
$\mathrm{VMO}^{\omega}_{\operatorname{far}}(\mathbb{R}^{n},Y)$ =$=\Big{\{}f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}\operatorname{\mathit{d}}(Q,0)>\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big{\}},$
@@
We continue by noting that the approximation $\mathrm{VMO}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}(\mathbb{R}^{n})}$ was already present in the classical work of Uchiyama [Uch1978]. Uchiyama’s construction does not easily translate to the case $\alpha\in(0,1)$ and thus in [GHWY21] a non-trivial approximation for $\mathrm{VMO}^{\alpha}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}^{\alpha}(\mathbb{R}^{n})}$ was provided that relies, more or less, on the existence of dyadic grids on the underlying space $\mathbb{R}^{n}$ (as does Uchiyama’s original approximation) and a careful construction using these grids. Such a construction, when $X$ is an arbitrary normed space, has no hope of working, simply because there is no grid available. Therefore, one more fact of our Theorem 1.6, in particular of its proof, is that it provides a conceptually easier approximation that uses only basic mollification and a carefully chosen truncation of the support. To make the connection with the mean oscillations and the pointwise conditions, we assume the following summability condition
@@
$\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\lesssim_{n,C_{\operatorname{db}}}[\omega]_{*}\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)},\qquad\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}\lesssim_{n,C_{\operatorname{db}}}[\omega]_{*}\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}.$ (8)
@@
The proof-idea of Theorem 1.9 goes back to Meyers [Mey1964], who obtained a version of the bounds (8) with the modulus $t^{\alpha},$ for $\alpha>0.$ The condition (7) appears new, but we recognize that it is a natural summability condition, after all. The identification (9) is completely new. Combining Theorems 1.6 and 1.9, we immediately obtain the following.
@@
1.5. Applications to compactness of the bi-commutator
Banach space valued operators and function spaces often appear in bi-parameter harmonic analysis. Then, it is maybe no surprise that the Banach valued Theorems (1.6), 1.10 and 1.9 have already found applications to bi-parameter harmonic analysis in the recent work [MarOik24]. We turn to the details.
We now consider $\mathbb{R}^{n}=\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}$ as a bi-parameter space, take non-degenerate CZOs $T_{i}$ on $\mathbb{R}^{n_{i}}$ and $b\in L^{1}_{\operatorname*{loc}}(\mathbb{R}^{n},\mathbb{C}).$ Then, the operator of interest, the bi-commutator, is
$[T_{2},[b,T_{1}]]=T_{2}[b,T_{1}]-[b,T_{1}]T_{2}=T_{2}bT_{1}-T_{2}T_{1}b-bT_{1}T_{2}+T_{1}bT_{2},$

### v4

@@ Abstract.
We introduce new vanishing subspaces of the homogeneous Hölder space $\dot{C}^{0,\omega}(X,Y)$ in the generality of a doubling modulus $\omega$ and normed spaces $X$ and $Y.$ For many couples $X,Y,$ we show these vanishing subspaces to completely characterize those Hölder functions that admit approximations, in the Hölder seminorm, by smooth, Lipschitz and boundedly supported functions. We present connections to bi-parameter harmonic analysis on the Euclidean space by providing applications to the compactness of the bi-commutator of two Calderón-Zygmund operators
@@ 1.1. Introduction
$|f|||_{C^{0,\omega}(X,Y)}=\|f\|_{\dot{C}^{0,\omega}(X,Y)}+\|f(0)\|_{Y}.$ (1)
Dropping the term $\|f(0)\|_{Y},$ the seminorm $\|\cdot\|_{\dot{C}^{0,\omega}(X,Y)}$ becomes a norm when the functions are identified with equivalence classes modulo additive constants. In the particular case $\omega(t)=t^{\alpha},$ $\alpha\in(0,1),$ we recover the Hölder spaces $\dot{C}^{0,\alpha}(X,Y).$ Moreover, provided $\omega$ is any reasonable modulus of continuity – at least $\omega(t)\to 0,$ as $t\to 0$ – the spaces $\dot{C}^{0,\omega}(X,Y)$ become subspaces of uniformly continuous functions.
@@
In the present article, we prove full characterizations of approximation with Hölder type rate of convergence taking the place of uniform convergence. We note that these two types of convergence do not follow from each other; see Appendix A for a detailed comparison with simple observations and examples. Our main result is best described as the recognition of the vanishing scales small, large and far as precisely the concept that allows a complete description of the closure of smooth, Lipschitz and boundedly supported functions with respect to the $\dot{C}^{0,\omega}$ seminorm. We notate
@@
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{small}}(X,Y)$ $=\big\{f:X\to Y\,:\,\lim_{\delta\to 0}\operatorname{osc}^{\omega}_{\delta}(f)=0\big\},\,$
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{large}}(X,Y)$ $=\big\{f:X\to Y\,:\,\lim_{\delta\to\infty}\operatorname{osc}^{\omega}_{\delta}(f)=0\big\},$
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{far}}(X,Y)$ $=\big\{f:X\to Y\,:\,\lim_{\delta\to\infty}\sup_{\min(\|x\|,\|y\|)>\delta}\operatorname{osc}^{\omega}_{(x,y)}(f)=0\big\}.$
@@
For each scale $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ define
@@
If $Y$ is a Banach space, then so are all of $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ as spaces of modulo additive constant equivalence classes; for full details, see Remark 2.1 below. From now on, whenever no confusion is possible, we start dropping the source and target spaces from the norms and simplify notation with $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|.$ In the same way, we write $\dot{C}^{0,\omega}(X,Y)=\dot{C}^{0,\omega}(X)=\dot{C}^{0,\omega},$ etc.
@@
where the doubling constant satisfies $C_{\operatorname{db}}>1.$ This class of moduli includes the Hölder moduli $\omega(t):=t^{\alpha},$ when $\alpha\in(0,1),$ but also other moduli that behave as $1/\log(1/t)$ close to the origin, which appear in the study of regularity for some partial differential equations, see e.g. Liao [LIAO2022]. Our results involve several assumptions on the moduli with different theorems requiring slightly different assumptions, see Table 1 below for a summary.
The classes $\dot{\operatorname{VC}}^{0,\alpha}_{\operatorname{small}}$ are also called the little Hölder spaces. These are important in the study of Lipschitz algebras in metric spaces, and we refer to the monograph by Weaver [Weaverbook, Chapters 4 and 8] for a background. See also the recent monograph by D. Mitrea, I. Mitrea, and M. Mitrea [Mitreabook, Chapter 3] for a detailed exposition of the Hölder spaces in the setting of Ahlfors regular sets and their connection with functions of vanishing mean oscillation.
@@ 1.3. Applications on the Euclidean space
Above we presented results that guarantee approximations provided we know pointwise vanishing type conditions. When $X=\mathbb{R}^{n},$ we will formulate these results in terms of mean oscillation type conditions, which are very convenient when studying commutators of Calderón-Zygmund operators (CZOs), for example. In the one parameter setting our results recover known components of the theory, but from a more general approach. In the bi-parameter context applications to the study of the compactness of commutators were recently presented in [MarOik24]. We describe all of this in full detail below, but let us begin by a motivation.
For many non-degenerate CZOs $T$ on $\mathbb{R}^{n},$ the symbol of the commutator $b\in L^{1}_{\operatorname*{loc}}(\mathbb{R}^{n},\mathbb{C})$ and $p,q\in(1,\infty),$ there holds that
@@
where the space $X^{p,q}(\mathbb{R}^{n})$ is determined as follows. Given $p,q\in(1,\infty),$ let $\alpha(p,q)$ and $r(p,q)$ be defined through $\alpha(p,q)/n=1/p-1/q$ and $1/q=1/r(p,q)+1/p.$ Then, $X^{p,q}(\mathbb{R}^{n})$ is the space of bounded mean oscillations $\mathrm{BMO}(\mathbb{R}^{n}),$ when $p=q;$ the (fractional BMO) space $\mathrm{BMO}^{\alpha(p,q)}(\mathbb{R}^{n}),$ when $p<q;$ and the Lebesgue space $L^{r(p,q)}(\mathbb{R}^{n})$ modulo additive constants $\dot{L}^{r(p,q)}(\mathbb{R}^{n}),$ when $p>q.$ (The definitions of the spaces $\mathrm{BMO},\mathrm{BMO}^{\alpha}$ can be read from Definition 1.7 below with the $\mbox{moduli}$" $t\mapsto t^{0}$ and $t\mapsto t^{\alpha},$ respectively.) With emphasis on the recognition of the correct spaces $X^{p,q}(\mathbb{R}^{n})$ to which the symbol belongs, the result (5) is due to Nehari [Nehari1957] ($n=1$) and Coifman, Rochberg, Weiss [CRW] ($n\geq 2$), when $q=p;$ due to Janson [Jan1978], when $q>p;$ and due to Hytönen [HyLpLq], when $q<p.$ We direct the reader to [HyLpLq] for a further discussion and precise definitions.
When $\alpha(p,q)>1,$ the space $\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n})$ consists of constant functions and the commutator is bounded iff it is the zero operator (hence compact, in particular). Thus, only those $p,q$ that result in $\alpha(p,q)\leq 1$ are interesting to study. When $\alpha(p,q)=1$, [GHWY21, Theorems 1.7. & 1.8.] show that the commutator is compact iff it is the zero operator. Then, for $p,q\in(1,\infty)$ in the non-trivial range $\alpha(p,q)<1,$ and for many non-degenerate CZOs $T$, there holds that
@@
The characterization (6) is due to Uchiyama [Uch1978], when $p=q$; due to Guo, He, Wu and Yang [GHWY21], when $q>p;$ and due to Hytönen, Li, Tao and Yang [HLTY2023], when $q<p.$ We recall that a result of Meyer’s [Mey1964] states that $\mathrm{BMO}^{\alpha(p,q)}(\mathbb{R}^{n})=\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n}),$ for $\alpha(p,q)\in(0,1).$ Thus, the following Theorem 1.6 provides in the case $\alpha(p,q)\in(0,1)$ a new characterization of the space $Y^{p,q}(\mathbb{R}^{n}).$
@@
To obtain the implications $\Rightarrow$" in (6) it is useful to have a description of $Y^{p,q}(\mathbb{R}^{n}),$ in the cases $q\geq p,$ in terms of vanishing mean oscillation (VMO) type criteria. Such VMO criteria (Definition 1.8, below) usually follow immediately from the existence of approximations, but to establish approximations from VMO type criteria is delicate. Recall that in the previous section we established the existence of approximations beginning from pointwise vanishing criteria $\dot{\operatorname{VC}}.$ Below, as Theorem 1.9, we provide the correspondence between these $\mathrm{VMO}$ and the $\dot{\operatorname{VC}}$ type criteria, which allows us, when $X=\mathbb{R}^{n},$ to reformulate all of our results, among them Theorem 1.6, in terms of mean oscillation conditions.
@@ 1.4. Formulation as mean oscillation conditions
We direct the reader to Section 1.4.1 below for a recap of Bochner integrability.
@@
$\mathrm{VMO}^{\omega}_{\operatorname{small}}(\mathbb{R}^{n},Y)$ $=\Big\{f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to 0}\sup_{\begin{subarray}{c}\ell(Q)=\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big\},$
$\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)$ $=\Big\{f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}\ell(Q)=\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big\},$
$\mathrm{VMO}^{\omega}_{\operatorname{far}}(\mathbb{R}^{n},Y)$ $=\Big\{f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}\operatorname{\mathit{d}}(Q,0)>\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big\},$
@@
The approximation $\mathrm{VMO}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}(\mathbb{R}^{n})}$ was already present in the work of Uchiyama [Uch1978]. Uchiyama’s construction relies on covering any bounded neighbourhood of the origin by disjoint cubes of equal shared side length $l>0$. The key property used is that there exists some $M>0$ such that given any cube of side length $\sim l$, it intersects at most $M$ many of these cubes. Uchiyama’s construction does not easily translate to the case $\alpha\in(0,1)$ and in [GHWY21] an approximation for $\mathrm{VMO}^{\alpha}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}^{\alpha}(\mathbb{R}^{n})}$ was provided that relies on the same covering property that Uchiyama’s original argument did. Clearly such a property fails, when $X$ is an arbitrary normed space. Therefore, one more fact of Theorem 1.6, in particular of its proof, is that it provides a conceptually easier approximation that uses only basic mollification along with a very carefully chosen truncation of the support. To make the connection with the mean oscillations and the pointwise conditions, we assume the following summability condition
@@
$\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\lesssim_{n}[\omega]_{*}\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)},\qquad\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}\lesssim_{n,C_{\operatorname{db}}}\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}.$ (8)
@@
The proof idea of Theorem 1.9 goes back to Meyers [Mey1964], who obtained a version of the bounds (8) with the modulus $t^{\alpha},$ for $\alpha>0.$ The condition (7) appears new, but we recognize that it is a natural summability condition, after all. The identification (9) is new. Combining Theorems 1.6 and 1.9, we immediately obtain the following.
@@
In the following table we gather the different background assumptions on modulus used in the theorems above.
Table 1. Conditions on the modulus in Theorems 1.2 through 1.10
Theorem 1.2 1.3 1.4 1.5 1.6 1.9 1.10
$\omega(0^{+})=0$ x x x x x - x
$\omega(\infty)=\infty$ x x x x x x x
$\omega$ is doubling x x x x x x x
$\lim_{t\to 0}t/\omega(t)=0$ x x x x x - x
$\omega]_{*}<\infty - - - - -$ x x
1.4.1. Bochner integrability
In the proof of Theorem 1.6, as well as in Theorem 3.12, we will make use of the Bochner integral to build up smooth approximations. We refer to [BenyaminiLindenstrauss, pp. 99–101] for an exposition on the integrability of Banach-valued maps, and to [HajJoh14, pp. 39–40] for Differentiation Under the Integral Sign results for the Bochner integral. If $\Omega\subset\mathbb{R}^{n}$ is Lebesgue measurable, a simple function $s:\Omega\to Y$ is a function of the form $s=\sum_{i=1}^{N}y_{i}\cdot\mathcal{X}_{E_{i}},$ for measurable subsets $E_{1},\ldots,E_{N}\subset\Omega$, and vectors $y_{1},\ldots,y_{N}\in Y.$ Then, we say that a function $f:\Omega\to Y$ is measurable if it is the a.e. pointwise limit of a sequence $\{s_{n}\}_{n}$ of simple funtions in $\Omega.$ Denoting by $\lambda$ the Lebesgue measure in $\mathbb{R}^{n}$, a simple function $s$ as above is Bochner-integrable in $\Omega$ whenever $\lambda(E_{i})<\infty$ for every $i,$ and the Bochner-integral of $s$ is defined by $\int_{\Omega}s(x)\,\mathrm{d}x=\sum_{i=1}^{N}y_{i}\lambda(E_{i}).$ And a measurable function $f:\Omega\to Y$ is Bochner-integrable if there exists a sequence of simple functions $\{s_{n}\}_{n}$ converging a.e. to $f$ in $\Omega$ so that
$\lim_{n,m\to\infty}\int_{\Omega}\|f_{n}(x)-f_{m}(x)\|_{Y}\,\mathrm{d}x=0.$
In such case, the Bochner integral of $f$ is defined by
$\int_{\Omega}f=\lim_{n}\int_{\Omega}s_{n}.$
Lastly, a function $f$ is said to be locally Bochner interable, provided that $f$ is Bochner integrable on each compact set $E\subset\mathbb{R}^{n}.$
1.5. Applications to the compactness of the bi-commutator
Theorems 1.6, 1.9 and 1.10 were partly motivated by applications to bi-parameter harmonic analysis that followed in the subsequent work of the second named author [MarOik24]. We turn to describing the details. We consider $\mathbb{R}^{n}=\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}$ as a bi-parameter space, take non-degenerate CZOs $T_{i}$ on $\mathbb{R}^{n_{i}}$ and $b\in L^{1}_{\operatorname*{loc}}(\mathbb{R}^{n},\mathbb{C})$, where $T_{i}$ acts on functions on $\mathbb{R}^{n}$ by acting only on the variable $x_{i}.$ Then, the bi-commutator $T_{2},[b,T_{1}]]$ is obtained by commuting the symbol first with $T_{1}$ and then with $T_{2}.$